### v1

@@ 3.1.3   Missing Data Imputation
The missing data imputation in transportation aims to reconstruct original data $X$ from missing or corrupted data $\hat{X}$ with conditional information $\phi$ such as known historical data, geometry configuration, and external factors ($\hat{X}=f(X,\phi)$). This process inherently involves data generation, as it creates new values to fill gaps in existing datasets, thereby enhancing the volume and quality of data available for analysis. In traditional imputation methods, missing data is often challenging to reconstruct due to the complex spatio-temporal dependencies in traffic data, which limit the accuracy and robustness of simple statistical methods. In recent years, DGMs have become powerful tools for handling the complexities of traffic data imputation. Their ability to model the underlying distribution of traffic data enables them to generate realistic and accurate reconstructions of missing entries across various scenarios. DGMs can also integrate a wide range of conditional information, including the spatio-temporal dependencies of traffic data, to enhance imputation performance (Zhang et al., 2024a; Cai et al., 2023; Yang et al., 2021; Zhang et al., 2021c; Peng et al., 2023; Shin et al., 2023; Duan et al., 2024; Chen et al., 2022b). Traditionally, missing data has been categorized into three types: Missing Completely at Random (MCAR), Missing at Random (MAR), and Missing Not at Random (MNAR) (Lin and Tsai, 2020; Hasan et al., 2021). However, the unique spatio-temporal characteristics of traffic data require a reevaluation of these categories. To address this, Chan et al. (2023) proposed a new classification specifically for transportation research, dividing missing data into Fiber Missing Data, Block Missing Data, and Random Missing Data. In this paper, we adopt this updated categorization to better address the specific challenges of traffic data imputation.
@@ 3.1.3.1   Fiber Missing Data
Fiber Missing Data occurs when there is a sudden, temporary failure of data-acquisition devices, leading to gaps in data collection that can be short-term or long-term. Many advanced models have been developed to address this issue effectively (Li et al., 2018; Zhang et al., 2021c, b; Huang and Chen, 2022; Yuan et al., 2022; Shen et al., 2022; Cai et al., 2023; Zhang et al., 2023b; Hou et al., 2023; Cai et al., 2023; Li et al., 2024c; Duan et al., 2024; Zhang et al., 2024a). For instance, Han et al. (2020) developed the Content-Aware GAN (CA-GAN) to handle missing traffic data in time series. In this method, traffic data were considered as tensors to capture their spatial-temporal correlations. The CA-GAN learns to generate realistic traffic patterns, effectively completing gaps in speed data. Similarly, Xu et al. (2021) introduced the Graph Aggregate GAN (GA-GAN), which combines the strengths of Graph Sample and Aggregate (GraphSAGE) with a GAN to impute missing traffic speed data. GraphSAGE aggregates information from neighboring nodes in the road network to capture spatial-temporal correlations, which are then used by a Wasserstein GAN (WGAN) to generate the missing data. The GA-GAN has been tested in both fiber and random missing data environments, demonstrating its versatility. Zhang et al. (2021b) proposed the Gated GAN (GaGAN) to address missing speed data on signalized roads. GaGAN incorporates attention mechanisms into graph convolutional operations and GRUs to learn spatio-temporal correlations within each signal cycle. This approach is effective for both fiber and random missing data scenarios. In another study, Liu et al. (2023) introduced PriSTI, a conditional diffusion framework for spatio-temporal data imputation. Specifically, the PriSTI uses a diffusion-based noise estimation module and a spatio-temporal feature extraction process to enhance its performance under high missing rates and diverse spatial configurations. It has been evaluated in environments with both fiber and random missing data, showing significant robustness. These models demonstrate the capabilities of DGMs in addressing the challenges posed by fiber missing data, ensuring the continuity and quality of traffic datasets even in the face of temporary data collection failures.
@@ 3.1.3.2   Block Missing Data
Block Missing Data occurs when there are no data-acquisition detectors in the area of interest, leading to complete loss in the dataset for the region. This type of missing data is particularly challenging to address because of the absence of any observations over an extended spatial and temporal range. Advanced DGMs have been developed to tackle this challenge, though most studies do not focus exclusively on block missing data. Instead, they often evaluate model performance under various conditions, including fiber and random missing data scenarios(Li et al., 2018; Zhang et al., 2021c; Yuan et al., 2022; Hou et al., 2023; Li et al., 2023b; Zhang et al., 2024a). For example, Yuan et al. (2022) enhanced the performance of GAN in traffic flow imputation by designing generative and center losses. This approach was evaluated across all three types of missing data environments: Fiber, Block, and Random Missing Data. The proposed method demonstrated significant improvements in imputing missing traffic data under these conditions. In another study, Li et al. (2018) combined a 3D Convolutional Neural Network with GAN (3wwDConvGAN) to address missing traffic data. Unlike traditional methods that may not fully exploit the spatial-temporal features of historical data, 3DConvGAN uses a fractional strided 3D CNN in both the generator and discriminator to enhance imputation performance. The model’s effectiveness was also tested in environments with Fiber, Block, and Random Missing Data, highlighting its versatility. Zhang et al. (2021c) introduced SA-GAIN, a Self-Attention Generative Adversarial Imputation Network, designed to impute missing traffic flow data. By incorporating a self-attention mechanism in GAN, SA-GAIN effectively captures important features across spatial and temporal dimensions. The model was tested in both fiber and block missing data scenarios, demonstrating robust performance. Hou et al. (2023) developed MissII, a model that combines spatial interaction models with GANs for traffic data imputation. This method first estimates traffic flow using theory-guided models that leverage environmental and social factors to capture complex traffic dynamics. The estimated traffic flow data are then used as real samples to guide the GAN’s training process to improve the imputation accuracy. These studies illustrate the effectiveness of DGMs in addressing block missing data, which presents unique challenges due to the complete absence of data in specific areas. By leveraging spatio-temporal modeling and integrating advanced neural network architectures, these models provide comprehensive solutions to ensure continuity and quality in traffic data even when entire sections of data are missing.
@@ 3.1.3.3   Random Missing Data
Random Missing Data occurs when the data-acquisition detectors fail unpredictably in the spatio-temporal domain, which results in gaps in data with little to no correlation among the missing entries. Due to its unpredictable nature, this scenario presents significant challenges and is commonly used to evaluate the performance of imputation models (Li et al., 2018; Chen et al., 2021a, 2022b; Wu et al., 2022a; Qin et al., 2021; Kazemi and Meidani, 2021; Tu et al., 2021; Zhang et al., 2021b; Xu et al., 2021; Shen et al., 2022; Yuan et al., 2022; Yang et al., 2022; Zhang et al., 2022a; Wang et al., 2023; Liu et al., 2023; Li et al., 2023b; Huang et al., 2023; Zhang et al., 2023b; Cai et al., 2023; Hou et al., 2023; Cai et al., 2023; Li et al., 2024c; Zheng et al., 2024; Duan et al., 2024; Zhang et al., 2024b, a). For example, Chen et al. (2019) developed a GAN-based method to impute missing traffic flow data by combining real and synthetic data. This approach utilizes a parallel data paradigm where the GAN generates synthetic data that are used alongside real data to train the imputation model, enhancing its ability to fill in gaps accurately. Boquet et al. (2019, 2020) used a VAE-based method for imputing missing traffic data, aimed at improving the accuracy of traffic forecasting systems impacted by sensor or system failures. The unsupervised approach learns the underlying data distribution from a latent space, resulting in improved post-imputation performance and effective data augmentation. To fully exploit the benefits of both VAE and GAN in data imputation, Shin et al. (2023) presented an Adversarial Autoencoder (AAE)-based model that leverages spatio-temporal feature extraction to address missing traffic data. AAE combines the principles of VAE and GAN to capture the complex dependencies within traffic data, providing robust imputation capabilities even in scenarios with high rates of missing data. Yang et al. (2021) introduced the Spatio-Temporal Learnable Bidirectional Attention GAN (ST-LBAGAN) for imputing missing traffic data. This model integrates U-Net and bidirectional attention mechanisms into a GAN-based framework to capture spatio-temporal dependencies for improving imputation performance. Additionally, Li et al. (2023a) proposed the Time Series Generative Adversarial Imputation Network (TGAIN) to tackle the challenge of imputing time series data. This model is designed to learn the multi-state distribution of missing time series data under conditional vector constraints. By employing a multiple imputation strategy, it effectively handles the uncertainty inherent in the imputation process. Recently, Zheng et al. (2024) used traffic domain knowledge and a denoising diffusion model to develop the Doubly Physics-Regularized Denoising Diffusion Model (DPRDDM) for recovering corrupted traffic speed data. This model demonstrates robustness in handling various types of noise, including Gaussian white noise, random corrupted noise, spatially correlated noise, temporally correlated noise, and a mixture of these noise types. These models illustrate the diverse applications and robust capabilities of DGMs in addressing the challenges posed by random missing data. By effectively capturing and utilizing spatio-temporal dependencies, these models ensure accurate and reliable data imputation even in the most unpredictable scenarios.
@@ 3.2.1   Agent-level Analysis
Agent-level analysis focuses on understanding and predicting the behaviors of individuals, such as vehicles, cyclists, or pedestrians. This detailed analysis is crucial for various applications like autonomous vehicle navigation, pedestrian safety systems, and dynamic routing. A significant portion of research in agent-level traffic analysis using DGMs primarily emphasizes learning the underlying distributions of traffic data (Roy et al., 2019; Zhao et al., 2020; Hegde et al., 2020; Zhou et al., 2021; Wang and He, 2021; Jagadish et al., 2021, 2022; Chen et al., 2022c; Hsu et al., 2023; Westny et al., 2024). This learning process is essential because it allows models to capture the intricate patterns and correlations within data that traditional models might miss. For example, Gómez-Huélamo et al. (2022) explored the use of attention mechanisms with GANs for vehicle trajectory prediction. The model generates feasible and realistic trajectories by considering both social interactions and the physical constraints of the road network. Similarly, Guo et al. (2023) developed the Map-Enhanced GAN (ME-GAN) for predicting autonomous vehicle trajectories. It leverages vehicle motion features, traffic state, and high-definition map constraints within the GAN framework to enhance prediction accuracy. Zhong et al. (2022) introduced the Spatial-Temporal Generative Model (STGM) to predict trajectories in autonomous driving environments. STGM combines a CVAE with dual spatial and temporal encoders to extract features from the driving environment. EquiDiff, proposed by Chen et al. (2023) is a conditional equivariant diffusion model that integrates a denoising diffusion probabilistic model with an SO(2)-equivariant transformer, effectively managing uncertainties in precise vehicle trajectory predictions.
@@
Another critical aspect of using DGMs in agent-level analysis is the ability to perform density estimation of outputs. This capability allows for generating multiple potential outcomes, offering a comprehensive probabilistic view that supports robust decision-making under uncertainty. Research has increasingly focused on this feature, enabling traffic managers and autonomous vehicle systems to adapt dynamically to varied and unpredictable conditions (Feng et al., 2019; Zhao et al., 2019; Cheng et al., 2020; Neumeier et al., 2021; Wang et al., 2020a; Eiffert et al., 2020; Rossi et al., 2021b; Choi et al., 2021b; Li et al., 2021; Rossi et al., 2021a; Oh and Peng, 2022; Xing et al., 2022; Jagadish et al., 2022; De Miguel et al., 2022; Gui et al., 2022; Yao et al., 2023; Tang et al., 2024; Lin et al., 2024). For example, Gupta et al. (2018) proposed the Social GAN to predict socially acceptable future trajectories for pedestrians in crowded scenes. The GAN model captures complex dynamics and social behaviors, generating a wide range of potential trajectories. Similarly, Kosaraju et al. (2019) introduced Social-BiGAT, a trajectory forecasting model that integrates Bicycle-GAN with Graph Attention Networks (GATs) and LSTMs. This model generates multimodal predictions of pedestrian trajectories by leveraging complex social interactions and physical context cues. In another study, Sun et al. (2021a) introduced a normalizing flow-based prediction model designed to model the exact probability distribution of future human trajectories. Kim et al. (2021) enhanced the accuracy of ego vehicle trajectory prediction using a driving style-based CVAE model. The proposed model integrates a DeepConvLSTM network for driving style recognition from in-vehicle sensor data, improving trajectory prediction. Wang et al. (2020c) presented a GAN-based framework for vehicle trajectory prediction that includes a novel auto-encoder social convolution and a recurrent social mechanism to capture both spatial and temporal interactions among vehicles. Expanding into air traffic, Wu et al. (2022b) proposed a GAN-based long-term 4D trajectory prediction method that generates multiple trajectories considering the spatiotemporal dynamics of flight paths. In a multi-agent context, Vishnu et al. (2023) incorporated traffic states into Transformer, GAN, and CVAE models to enhance multi-agent trajectory predictions. These models generate plausible trajectory outcomes that are contextually relevant and highly predictive. Salzmann et al. (2020) developed Trajectron++, an advanced VAE-based model that forecasts the trajectories of multiple agents in dynamic environments. Trajectron++ integrates a CVAE with Gaussian Mixture Models (GMMs) within a graph-based recurrent neural network framework. This structure allows Trajectron++ to generate multiple potential trajectories in a probabilistic manner, providing a comprehensive representation of possible future scenarios. Li et al. (2019a) proposed the Conditional Generative Neural System (CGNS), which combines the CVAE with GAN to generate feasible, realistic, and diverse future trajectories for multiple agents by leveraging both static context and dynamic interactions. Bhattacharyya et al. (2019) introduced Conditional Flow VAE (CF-VAE), which uses conditional normalizing flows to improve the forecast of future states by capturing complex, multi-modal conditional distributions. Li et al. (2023c) developed a framework combining CVAE and conditional diffusion models for multi-modal vehicle trajectory prediction. This approach addresses the challenge of predicting highly uncertain future vehicle trajectories in urban environments by first generating trajectories with CVAE and refining them using a diffusion model. Similarly, Liu et al. (2024) aimed to predict the distribution of endpoints of multi-agent trajectories using denoising diffusion and Transformer models, capturing both the spatio-temporal dynamics and the intrinsic intent of vehicles to enhance overall prediction quality. These examples illustrate the diverse and robust capabilities of DGMs in agent-level analysis, enabling precise and reliable predictions that are critical for the development and operation of intelligent transportation systems.

### v3

@@ 3.1.3   Missing Data Imputation
The missing data imputation in transportation aims to reconstruct original data $X$ from missing or corrupted data $\hat{X}$ with conditional information $\phi$ such as known historical data, geometry configuration, and external factors ($\hat{X}=f(X,\phi)$). This process inherently involves data generation, as it creates new values to fill gaps in existing datasets, thereby enhancing the volume and quality of data available for analysis. In traditional imputation methods, missing data is often challenging to reconstruct due to the complex spatio-temporal dependencies in traffic data, which limit the accuracy and robustness of simple statistical methods. In recent years, DGMs have become powerful tools for handling the complexities of traffic data imputation. Their ability to model the underlying distribution of traffic data enables them to generate realistic and accurate reconstructions of missing entries across various scenarios. DGMs can also integrate a wide range of conditional information, including the spatio-temporal dependencies of traffic data, to enhance imputation performance (Zhang et al., 2024a; Cai et al., 2023; Yang et al., 2021; Zhang et al., 2021c; Peng et al., 2023; Shin et al., 2023; Duan et al., 2024; Chen et al., 2022b). Traditionally, missing data has been categorized into three types: Missing Completely at Random (MCAR), Missing at Random (MAR), and Missing Not at Random (MNAR) (Lin and Tsai, 2020; Hasan et al., 2021). However, the unique spatio-temporal characteristics of traffic data require a reevaluation of these categories. To address this, Chan et al. (2023) proposed a new classification specifically for transportation research, dividing missing data into Fiber Missing Data, Block Missing Data, and Random Missing Data. In this paper, we adopt this updated categorization to better address the specific challenges of traffic data imputation. Table 4 summarizes recent studies and classifies them according to missing type, imputation task, model, dataset, missing rate, and evaluation metrics.
Table 4: A summary of recent studies in missing data imputation using DGM
Missing Type Author Imputation Task Model Dataset Missing Rate Evaluation Metrics
Fiber Han et al. (2020) Speed CA-GAN PEMS 10–100% MRE, Time Cost Comparison
Xu et al. (2021) Speed GA-GAN PEMS-BAY, Seattle Dataset 10–70% MAE, RMSE, MAPE, Residual Analysis
Zhang et al. (2021b) Lane-level Speed GaGAN Hangzhou Signalized Road Data 20–100% MAE, RMSE, MAPE, Pearson’s Correlation Coefficient
Liu et al. (2023) Speed PriSTI (Diffusion-based) AQI-36, METR-LA, PEMS-BAY 10–90% MAE, MSE, CRPS
Block Li et al. (2018) Flow 3DConvGAN Beijing Taxi Dataset 20–80% RSE
Zhang et al. (2021c) Volume SA-GAIN (GAN-based) Seattle I-5 Highway 10–80% MAE, MMD, RMSE
Yuan et al. (2022) Speed Passenger Flow STGAN Beijing Road Dataset Beijing Subway Dataset 20–80% RMSE, NMAE, MAE
Hou et al. (2023) Flow MissII (GAN-based) Beijing Taxi Dataset 20–60% MAE, RMSE, CosineSim
Random Chen et al. (2019) Flow GAN PEMS 30–80% MAE, RMSE, MRE
Boquet et al. (2019) Speed VAE PEMS 10–40% RMSE, MAPE
Shin et al. (2023) Speed AAE Korean ITS Data, PEMSD7 10–50% RMSE, MAPE
Yang et al. (2021) In&Outflow ST-LBAGAN Beijing Taxi, NYC Bike Datasets 10–60% RMSE, MAE
Li et al. (2023a) Volume Speed TGAIN (GAN-based) I90 Dataset, Changchun Speed Dataset 10–90% RMSE, MAPE
Zheng et al. (2024) Speed DPRDDM (Diffusion-based) Zen & Beijing Traffic Data 10–30% RMSE, MAPE
@@ 3.1.3.1   Fiber Missing Data
Fiber Missing Data occurs when there is a sudden, temporary failure of data-acquisition devices, leading to gaps in data collection that can be short-term or long-term. Many advanced DGMs have been developed to address this issue effectively (Li et al., 2018; Zhang et al., 2021c, b; Huang and Chen, 2022; Yuan et al., 2022; Shen et al., 2022; Cai et al., 2023; Zhang et al., 2023b; Hou et al., 2023; Cai et al., 2023; Li et al., 2024c; Duan et al., 2024; Zhang et al., 2024a). For instance, Han et al. (2020) developed the Content-Aware GAN (CA-GAN) to handle missing traffic data in time series. In this method, traffic data were modeled as tensors to capture the inherent spatial-temporal correlations. The CA-GAN learns the underlying traffic distribution and generates realistic traffic patterns, effectively completing gaps in speed data, particularly in cases with consecutive data losses. Similarly, Xu et al. (2021) introduced the Graph Aggregate GAN (GA-GAN), which combines the strengths of Graph Sample and Aggregate (GraphSAGE) with a GAN to impute missing traffic speed data. In this approach, GraphSAGE aggregates information from neighboring nodes in the road network to capture spatial-temporal correlations, which are then used by a Wasserstein GAN (WGAN) to generate the missing data. The GA-GAN has been tested in both fiber and random missing data environments, demonstrating its versatility. Zhang et al. (2021b) proposed the Gated GAN (GaGAN) to address missing speed data on signalized roads. The GaGAN integrates attention mechanisms within graph convolutional operations to capture spatial correlations and enhances GRUs with Self-Attention (SA-GRU) to learn temporal dependencies across signal cycles. This approach is effective for both fiber and random missing data scenarios, ensuring that the imputed values retain the inherent spatio-temporal patterns observed in lane-level traffic measurements. In another study, Liu et al. (2023) introduced PriSTI, a conditional diffusion framework for spatio-temporal data imputation. Specifically, the PriSTI uses a diffusion-based noise estimation module and a spatio-temporal feature extraction process to enhance its performance under high missing rates and diverse spatial configurations. It has been evaluated in environments with both fiber and random missing data, showing significant robustness. These models demonstrate the capabilities of DGMs in addressing the challenges posed by fiber missing data, ensuring the continuity and quality of traffic datasets even in the face of temporary data collection failures.
@@ 3.1.3.2   Block Missing Data
Block Missing Data occurs when there are no data-acquisition detectors in the area of interest, leading to complete loss in the dataset for the region. This type of missing data is particularly challenging to address because of the absence of any observations over an extended spatial and temporal range. Advanced DGMs have been developed to tackle this challenge, though most studies do not focus exclusively on block missing data. Instead, they often evaluate model performance under various conditions, including fiber and random missing data scenarios (Li et al., 2018; Zhang et al., 2021c; Yuan et al., 2022; Hou et al., 2023; Li et al., 2023b; Zhang et al., 2024a). For example, Li et al. (2018) combined a 3D Convolutional Neural Network with GAN (3DConvGAN) to address missing traffic data. Unlike traditional methods that may not fully exploit the spatial-temporal features of historical data, 3DConvGAN uses a fractional strided 3D CNN in both the generator and discriminator to enhance imputation performance. The model’s effectiveness was also tested in environments with Fiber, Block, and Random Missing Data, highlighting its versatility. In another study, Zhang et al. (2021c) introduced SA-GAIN, a Self-Attention Generative Adversarial Imputation Network, designed to impute missing traffic flow data. By incorporating a self-attention mechanism in GAN, SA-GAIN effectively captures important features across spatial and temporal dimensions. The model was tested in both fiber and block missing data scenarios, demonstrating robust performance. In Yuan et al. (2022), the authors enhanced the performance of GAN in traffic flow imputation by designing generative and center losses. These losses enable the model to more accurately reconstruct missing data while preserving local spatio-temporal correlations. This approach was evaluated across all three types of missing data environments: Fiber, Block, and Random Missing Data. The proposed method demonstrated significant improvements in imputing missing traffic data under these conditions. Hou et al. (2023) developed MissII, a novel theory-guided deep learning framework for traffic data imputation. The approach first estimates traffic flow between points of interest using mobility models that leverage environmental and social factors to capture complex traffic dynamics. The estimated traffic flow data are then used as real samples to guide the GAN’s training process to improve the imputation accuracy. These studies illustrate the effectiveness of DGMs in addressing block missing data, which presents unique challenges due to the complete absence of data in specific areas. By leveraging spatio-temporal modeling and integrating advanced neural network architectures, these models provide comprehensive solutions to ensure continuity and quality in traffic data even when entire sections of data are missing.
@@ 3.1.3.3   Random Missing Data
Random Missing Data occurs when the data-acquisition detectors fail unpredictably in the spatio-temporal domain, which results in gaps in data with little to no correlation among the missing entries. Due to its unpredictable nature, this scenario presents significant challenges and is commonly used to evaluate the performance of imputation models (Li et al., 2018; Chen et al., 2021a, 2022b; Wu et al., 2022a; Qin et al., 2021; Kazemi and Meidani, 2021; Tu et al., 2021; Zhang et al., 2021b; Xu et al., 2021; Shen et al., 2022; Yuan et al., 2022; Yang et al., 2022; Zhang et al., 2022a; Wang et al., 2023; Liu et al., 2023; Li et al., 2023b; Huang et al., 2023; Zhang et al., 2023b; Cai et al., 2023; Hou et al., 2023; Cai et al., 2023; Li et al., 2024c; Zheng et al., 2024; Duan et al., 2024; Zhang et al., 2024b, a). For example, Chen et al. (2019) developed a GAN-based method to impute missing traffic flow data by combining real and synthetic data. This approach utilizes a parallel data paradigm where the GAN generates synthetic data that are used alongside real data to train the imputation model, enhancing its ability to fill in gaps accurately. Boquet et al. (2019, 2020) used a VAE-based method for imputing missing traffic data, aimed at improving the accuracy of traffic forecasting systems impacted by sensor or system failures. The unsupervised approach learns the underlying data distribution from a latent space, resulting in improved post-imputation performance and effective data augmentation. To fully exploit the benefits of both VAE and GAN in data imputation, Shin et al. (2023) presented an Adversarial Autoencoder (AAE)-based model that leverages spatio-temporal feature extraction to address missing traffic data. AAE combines the principles of VAE and GAN to capture the complex dependencies within traffic data, providing robust imputation capabilities even in scenarios with high rates of missing data. Yang et al. (2021) introduced the Spatio-Temporal Learnable Bidirectional Attention GAN (ST-LBAGAN) that combines a U-Net generator with bidirectional attention to capture spatio-temporal dependencies for missing traffic data imputation. The model uses multi-channel inputs with a mask to focus on missing regions and is trained with a composite loss to ensure realistic outputs, achieving high accuracy even at high missing rates. Additionally, Li et al. (2023a) proposed the Time Series Generative Adversarial Imputation Network (TGAIN) to tackle the challenge of imputing time series data. This model is designed to learn the multi-state distribution of missing time series data under conditional vector constraints. By employing a multiple imputation strategy, it effectively handles the uncertainty inherent in the imputation process. Recently, Zheng et al. (2024) used traffic domain knowledge and a denoising diffusion model to develop the Doubly Physics-Regularized Denoising Diffusion Model (DPRDDM) for recovering corrupted traffic speed data. This model demonstrates robustness in handling various types of noise, including Gaussian white noise, random corrupted noise, spatially correlated noise, temporally correlated noise, and a mixture of these noise types. The current studies in random missing scenario illustrate the diverse applications and robust capabilities of DGMs in addressing the challenges posed by random missing data. By effectively capturing and utilizing spatio-temporal dependencies, these models ensure accurate and reliable data imputation even in the most unpredictable scenarios.
DGMs show promising solutions for handling missing data in traffic systems; however, some key challenges still persist at this stage. A primary concern is limited transferability: models tend to perform well on their training datasets but struggle to generalize across diverse traffic conditions and sensor networks, limiting their practical applicability. Additionally, high missing rates and irregular patterns substantially affect imputation quality, compounded by the difficulty of capturing complex spatiotemporal dependencies in multidimensional traffic data. Furthermore, the lack of standardized evaluation protocols—such as consistent missing rate ranges, evaluation metrics, and unified datasets—makes it difficult to compare and validate different approaches effectively. These limitations highlight the need for further research to improve the robustness and adaptability of DGMs in addressing missing data in traffic systems.
@@ 3.2.1   Agent-level Analysis
Agent-level analysis focuses on understanding and predicting the behaviors of individuals, such as vehicles, cyclists, or pedestrians. This detailed analysis is crucial for various applications like autonomous vehicle navigation, pedestrian safety systems, and dynamic routing. DGMs are particularly critical in this context since they can learn complex agent behavior patterns directly from data while inherently accounting for uncertainty in the estimation or prediction process. In other words, DGMs can reflect the inherent variability of real-world actions and results in more robust and reliable analysis. Table 5 summarizes recent studies, categorizing them by model, observation/prediction length, sampling time, prediction samples, dataset, and evaluation metrics.
Table 5: A summary of recent studies in agent-level analysis using DGM
Author Model Observation / Prediction Length Sampling Time Prediction Samples Dataset Evaluation Metrics
Li et al. (2019a) GAN + cVAE 2s/5s 3.2s/4.8s 0.5s 0.4s 1 INTERACTION ETH, UCY, Stanford Drone Dataset ADE, FDE
Kim et al. (2021) cVAE 1s/1s 0.1s 1 CarMaker HILS RMSE, MAE
Wang et al. (2020c) TS-GAN 3s/1-5s 0.1s 1 NGSIM RMSE
Gómez-Huélamo et al. (2022) GAN 2s/3s 0.1s 1 Argoverse ADE, FDE
Chen et al. (2023) EquiDiff 3s/1-5s 0.2s 1 NGSIM RMSE
Liu et al. (2024) Diffusion 3s/1-5s 0.2s 1 NGSIM RMSE
Gupta et al. (2018) SocialGAN 3.2s/4.8s 0.4s 1-100 ETH, UCY minADEk, minFDEk
Kosaraju et al. (2019) Social-BiGAT (GAN-based) 3.2s/4.8s 0.4s 1-20 ETH, UCY minADEk, minFDEk
Sun et al. (2021a) Flow 3.2s/0.4-4.8s 0.4s 1-20 ETH, UCY, Stanford Drone Dataset minADEk, minFDEk, NLL
Salzmann et al. (2020) Trajectron++ (VAE-based) 3.2s/4.8s 2s/1-4s 0.4s 0.5s 20 ETH, UCY nuScenes minADEk, minFDEk, KDENLLk
Vishnu et al. (2023) TS-GAN, TS-CVAE 3.2s/4.8s 0.4s 1-20 EyeonTraffic, INTERACTION minADEk, minFDEk, KDENLLk
Li et al. (2023c) cVAE + Diffusion 2s/6s 2s/3s 0.5s 0.1s 1-15 nuScenes Argoverse minADEk, minFDEk
@@
Notably, a significant portion of research in agent-level traffic analysis using DGMs primarily emphasizes learning the underlying distributions of traffic data and produces single outcome(Roy et al., 2019; Zhao et al., 2020; Hegde et al., 2020; Zhou et al., 2021; Wang and He, 2021; Jagadish et al., 2021, 2022; Chen et al., 2022c; Hsu et al., 2023; Westny et al., 2024). This learning process is essential because it allows models to capture the intricate patterns and correlations within data that traditional models might miss. For example, Li et al. (2019a) proposed the Conditional Generative Neural System (CGNS), which combines the cVAE with GAN to generate feasible, realistic, and diverse future trajectories for multiple agents by leveraging both static context and dynamic interactions. Kim et al. (2021) improved the accuracy of ego vehicle trajectory prediction using a driving style-based cVAE. Their model integrates a DeepConvLSTM network to recognize driving styles from in-vehicle sensor data. This approach not only predicts a single trajectory but also estimates a probability distribution over future trajectories, effectively capturing the inherent uncertainty in driving behavior. Additionally, Wang et al. (2020c) presented a GAN-based framework TS-GAN for vehicle trajectory prediction that uses multi-vehicle collaborative learning. Their method combines an auto-encoder social convolution module to capture spatial interactions and a recurrent social mechanism to model temporal relationships among surrounding vehicles. These fused features feed into a conditional GAN that generates a multi-modal probability distribution over future trajectories for a target vehicle. Similarly, Gómez-Huélamo et al. (2022) explored the use of attention mechanisms with GANs for vehicle trajectory prediction. The model generates feasible and realistic trajectories by considering both social interactions and the physical constraints of the road network. Recently, Chen et al. (2023) introduced EquiDiff, a conditional equivariant diffusion model that combines a denoising diffusion probabilistic model with an SO(2)-equivariant transformer to effectively manage uncertainties in vehicle trajectory predictions Liu et al. (2024) aimed to predict the distribution of endpoints of multi-agent trajectories using denoising diffusion and Transformer models, capturing both the spatio-temporal dynamics and the intrinsic intent of vehicles to enhance overall prediction quality.
@@
Another critical aspect of using DGMs in agent-level analysis is the ability to perform density estimation of outputs. This capability allows for generating multiple potential outcomes, offering a comprehensive probabilistic view that supports robust decision-making under uncertainty. Research has increasingly focused on this feature, enabling traffic managers and autonomous vehicle systems to adapt dynamically to varied and unpredictable conditions (Feng et al., 2019; Zhao et al., 2019; Bhattacharyya et al., 2019; Cheng et al., 2020; Neumeier et al., 2021; Wang et al., 2020a; Eiffert et al., 2020; Agarwal et al., 2020; Dendorfer et al., 2021; Rossi et al., 2021b; Choi et al., 2021b; Li et al., 2021; Rossi et al., 2021a; Oh and Peng, 2022; Wu et al., 2022b; Zhong et al., 2022; Guo et al., 2023; Xing et al., 2022; Jagadish et al., 2022; De Miguel et al., 2022; Gui et al., 2022; Yao et al., 2023; Rempe et al., 2023; Tang et al., 2024; Yang et al., 2024). For example, Gupta et al. (2018) proposed the Social GAN to predict socially acceptable future trajectories for pedestrians in crowded scenes. The GAN model captures complex dynamics and social behaviors, generating a wide range of potential trajectories. Extending this idea, Kosaraju et al. (2019) introduced Social-BiGAT, a trajectory forecasting model that integrates Bicycle-GAN with Graph Attention Networks (GATs) and LSTMs. This model generates multimodal predictions of pedestrian trajectories by leveraging complex social interactions and physical context cues. Other studies have adopted different strategies to model uncertainty. Sun et al. (2021a) introduced a normalizing flow-based prediction model designed to model the exact probability distribution of future human trajectories. In a multi-agent context, Salzmann et al. (2020) developed Trajectron++, an advanced VAE-based model that forecasts the trajectories of multiple agents in dynamic environments. Trajectron++ integrates a CVAE with Gaussian Mixture Models (GMMs) within a graph-based recurrent neural network framework. This structure allows Trajectron++ to generate multiple potential trajectories in a probabilistic manner, providing a comprehensive representation of possible future scenarios. Vishnu et al. (2023) incorporated traffic states into Transformer, GAN, and CVAE models to enhance multi-agent trajectory predictions. These models generate diverse plausible trajectory outcomes that are contextually relevant and highly predictive. Recent work has further explored hybrid approaches. Li et al. (2023c) developed a framework combining cVAE and conditional diffusion models for multi-modal vehicle trajectory prediction. This approach addresses the challenge of predicting highly uncertain future vehicle trajectories in urban environments by first generating trajectories with cVAE and refining them using a diffusion model.
These examples underscore the diverse and robust capabilities of DGMs in agent-level analysis, which are crucial for intelligent transportation systems. However, as models incorporate richer contextual data and accommodate a variable number of agents, their computational complexity increases, posing challenges for real-time scalability. Moreover, effectively integrating complex dynamics and heterogeneous contextual cues remains a significant difficulty.